A Go-ped traveled 9 miles in 24 minutes.
How long would it take to go 27 miles?
Thanks
: A Go-ped traveled 9 miles in 24 minutes.
How long would it take to go 27 miles?
Thanks

Answer by scott8148(2761)

27 miles is 3 times as far as 9 miles (27/9=3), so it takes 3 times as long
__ 24*3=72

let x=time for 27 miles __ the distance is PROPORTIONAL to the time

9/24=27/x __ "cross" multiplying __ 9x=27*24 __ dividing by 9 __ x=3*24 __ x=72
Question 147656: A Go-ped traveled 9 miles in 24 minutes.
How long would it take to go 27 miles?
Thanks
: A Go-ped traveled 9 miles in 24 minutes.
How long would it take to go 27 miles?
Thanks

Answer by Fombitz(1756)

Set up a ratio.
9 miles is to 24 minutes as 27 miles is to X minutes.
Mathematically it would look like this,
9/24=27/Y
Y=27*(24/9)
Y=72
It would take 72 minutes to go 24 miles.